Circular No. 6368

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1987A IN THE LARGE MAGELLANIC CLOUD P. Garnavich and R. Kirshner report, on behalf of the Supernova IN-

tensive Study (SINS) collaboration: “Hubble Space Telescope images of SN 1987A show that the [N II] 658.4-nm flux from a segment of the inner circumstellar ring increased by 15 percent between 1994 Oct. and 1996 Feb. The ring segment is centered at p.a. 206^{o} and subtends ∼ 35^{o}. The remainder of the ring has faded by 5 to 10 percent in [N II]. No unusual activity is seen the [O III] 500.7-nm emission, and the entire ring faded by an average of 10 percent. The flux increase could be the first indication of the predicted interaction between the supernova ejecta and the circumstellar gas. Observations at radio and x-ray wavelengths may show considerable changes if the interaction has begun. However, since the optical activity is detected only in [N II] and is on the south side (thought to be the far side) of the ring, it is unlikely to be the onset of the collision. Models by Lundqvist and Fransson (1996, Ap.J., in press) imply that [N II] emission would continue to increase beyond day 2000 if the ring segment consisted of very-low-density gas.”

GRO J1744–28 T. Dotani, Y. Ueda, F. Nagase, and H. Inoue, Institute of Space and

Astronautical Science; and C. Kouveliotou, Universities Space Research Association and Marshall Space Flight Center, NASA, communicate: “Further analysis of ASCA data of GRO J1744–28 obtained between Feb. 26.42 and 27.09 UT yields the source position at α = 17^{h}44^{m}31^{s}.5, δ = −28^{o}44′26′′ (equinox 2000.0), with a 90-percent-confidence error radius of 1′. The source position was also cross-checked with reference to the Einstein source 2E 1743.1–2842, which was detected in the same field of view. The result is fully consistent with the above coordinates, taking into account the IPC error radius of 51′′ for 2E 1743.1–2842. The ASCA error circle overlaps the XTE error region (IAUC 6309), but the VLA radio source (IAUC 6323) lies slightly outside the ASCA error circle.”

COMET 65P/GUNN Visual m_{1} estimates: Feb. 19.51 UT, 13.6 (A. Hale, Cloudcroft, NM,

0.41-m reflector); 25.10, 13.6 (K. Sárneczky, Ráktanya, Hungary, 0.44m reflector); Mar. 17.48, 13.3 (Hale); 26.45, 13.4 (Hale); 28.07, 13.4 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector).

1996 April 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green